2nd EU-JA Fusion Neutron Source workshop

04 & 05 October 2021

The 2nd Fusion Neutron Source workshop (FNS WS #2) was held remotely on October 04th and 05th.

The workshop provided a general overview of the FNS collaboration activities that are carried out in Japan and Europe within the IFMIF/EVEDA, DONES and A-FNS projects.

Additional common topics of interest have been identified. They will be further discussed in the future, in order to complement the current dense work programme already agreed among all.

Although short, the workshop aroused interest for deep technical discussions and we are looking forward to meeting in person to continue these discussions in the forthcoming workshop.

View of FNS WS #2 remote participants

Weekly LIPAc highlights (8/November-12/November)

Injector

Deuteron beam extraction for injector conditioning was carried out until 26 October and reached 30 mA @ 10% duty cycle. On 27 October proton beam extraction was started and reached quickly the current and duty level required to resume beam operation (total extracted current 20mA, duty 3%). 

RF system

One of the new Driver Anode Box (chain 1B) had an issue and was recovered by replacing it back to the old one (checks are on-going). Anode Boxes in the remaining chains will be replaced progressively in the next short term maintenances.

RF conditioning started on 28 October: in a half day, 1% duty (1ms, 10Hz) was reached. The system could very quickly ramp-up to 10% duty cycle, confirming the much improved stability of the overall RF system. Duty cycle reached 10% at 2ms, 50Hz. During last weekend at ~17% duty cycle there was an important RFQ vacuum interlock event and associated RFQ coupler temperature increase. The system recovered and was reconditioned at low duty cycle with no issue. A maximum of 10% duty cycle was kept with varying pulse length and repetition rate. Upon request by INFN, field flatness measurement were performed at several duty cycle (with fix pulse duration of 4ms) before going to the next step increasing the duty cycle up to 15%. Until winter break the plan is to have beam operation in day time and keep conditioning during the night time and weekend.

Beam Transport/BD

Before resuming beam operation, there was a very intense period of remote sessions with CIEMAT to solve some last minute issue and complete few remaining tasks. Power Supply (PS) OPI was modified to allow the use of the PS HPU04 which is perfectly functional but does not provide digital current readback (analog output are operative and processed through the OPI). 

Diagnostics

In the remote sessions with CIEMAT the tasks for the cRIO SW update, the noise issues of BD accelerometer and the water temperature measurement at the D-slit were completed. BPM clock board was recovered by applying SW update.

Ancillaries/vacuum and cooling

Just before resuming the beam one vacuum gate valve (isolating a cryopump in the HEBT vacuum sector III) was blocked half open. Despite all the efforts it could not be recovered. A number of countermeasures have been taken to operate safely the vacuum system during the beam and the intervention on the valve for the next winter break is in preparation.

Improvements of the secondary cooling system (outdoor large water tank area) restarted to install protector against drop of ice and frozen snow.

The large Cryoplant 2nd storage buffer recently installed has been connected, conditioned and filled with some pure helium.

Control System

MPS test before beam operation was longer than planned and it was necessary to have CIEMAT to get all MPS signals for the DPlate diagnostic cleared. The MPS for the HEBT/BD was completed on 11 November in the morning.

Beam operation

A rehearsal for beam operation was done on 11 November and beam operation started on November 12.

Operations Supervisor and sub-systems operators observing traces of first beams in CCR
Remote participants observing traces of first beam on shared OPI

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Weekly LIPAc highlights (11/October-15/October)

Injector

The maintenance of injector in the vault was completed. Accelerator column alignment was checked and found to be in line with last year checks. Boron Nitride disks were found to be strongly damaged after half year operation: they have been replaced by new ones (after baking). The cause of damage (surface eroded and colored in black) needs to be identified. Spares Boron Nitride disks are available on-site.

The PPS test was performed on Tuesday and HV up to 100kV were applied with no issues. D+ plasma conditioning was resumed on Wednesday this week.

Radio Frequency (RF) System and RFQ cavity

The last activities before resume RF-RFQ modules were completed, namely the installation of the re-calibrated VSWR cards into the PSYS units and the installation of the Anode Boxes with new PCB circuits into RFQ module 1 and 4B. The start-up process of RFQ-RF modules started on Wednesday.

An accidental spill of water occurred from a 50 kW dummy load during a series of planned tests on the SSPA-A. To check the SSPA functionalities after the accident, tests procedures were prepared and tests carried out: the SSPA-A functionality is unaltered, with a nominal output power extracted from the SSPA modules. A damage was found in an electronic board of the peripheral control unit and spare procurement started.

On the RFQ manifolds in the vault, the check of water flow rates on the hundreds of individual channels equipped with flow regulators was completed and results confirmed that no immediate actions is needed.

Beam Transport

The checks and calibration of the HEBT magnet PS is almost completed.

The mechanical check of the MEBT bunchers has been completed.

Functional checks of mass flow controllers for HEBT vacuum pressure control is on-going with CIEMAT remote support.

Ancillaries

A new radiation resistant RGA and a new rupture disk were installed in vacuum sector IV of the HEBT.

All cryopumps for RFQ and HEBT have been successfully restarted.

Diagnostics

Newly calibrated HEBT FPMs (Fluorescence Profile Monitors) were installed. Functional check will be done though the remote session with CIEMAT.

Checks are ongoing on BD accelerometer cables with CIEMAT remote support.

Control system.

The integration of common chiller LCS into the CCS is on-going. The communication line between BD skid and HEBT LCS has been renewed.

In the current Integrated LIPAc schedule, it is planned to resume beam operation early November.

HEBT sector IV, new RGA radiation resistant + new rupture disk

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Weekly LIPAc highlights (27/Sepember-1/October)

As you may remember the LIPAc beam operation ended at the end of July.

Starting from August and until end of September there are major maintenance interventions in LIPAc, including regulatory inspections (on electrical system, cooling systems, HVAC, crane in the vault etc.) compulsory by law.

Injector

The new Emittance Meter Unit delivered by F4E has been installed in the vault, cleaned and the compatibility with the ancillaries and mechanical connections confirmed.

The main maintenance on the injector has just started and will continue until early October.

Injector conditioning is planned to resume by mid of October.

Radio Frequency (RF) System and RFQ cavity

The full maintenance of HVPS and RF-RFQ modules was performed with the support of NAT subcontractor. The RF team is currently working on upgrades of the RFQ RF modules to improve their functionality and stability in operation: upgraded PCB cards for the anode boxes have been validated and are being installed; PSYS VSWR cards to fit the new predrivers are under calibration; AC filters have been installed inside the PLC for noise reduction purposes.

By mid-October the RFQ RF modules will resume operation and target end of October to start injection in the RFQ. RF team will then need one week to perform three activities that needs RF injection in the cavity and then the RF will be ready for beam operation.

Beam Transport

Maintenance, checks and calibration of MEBT and MEL magnets is almost completed. HEBT magnets calibration is planned next week.

A new radiation resistance RGA and a new rupture disk will be installed this week in HEBT vacuum sector IV. The calibration and cable routing for the RGA are completed.

Ancillaries

A part the standard maintenance, several interventions were performed in the cooling skids (active carbons replacement and UPS replacement in the RF skid, valves inspections and recovery of functionality). Some pending interventions on the skids requires specific skills and a dedicated contract is being prepared by QST targeting the execution at the end of the year.

Also the newly commissioned Beam Dump skid underwent a thorough maintenance in September.

Diagnostics

Remote sessions with CIEMAT support have resumed to complete the commissioning of the HEBT and Beam Dump instrumentation (Faraday Cup, Fluorescence Profile Monitors, SEM grids motion control), debug HEBT PLC and complete the commissioning of the HEBT vacuum machine state. The pending hardware and cabling installation work was mostly executed onsite in August and remote sessions have been resumed in September

A VME control issues has to be solved to allow completing all the checks on Current Transformers and SEM grids data acquisition. CEA is supporting remotely.

A BPM in the MEBT is still not fully operational: the plan to remove the upper half of the quadrupole (MEBT BPM are embedded in the magnets) to access the BPM has been postponed to avoid the risk of the intervention.

Cryoplant

In September a large Helium buffer tank was installed using large cranes (see pictures below).

André Duarte was on-site for a month in July-August (first mission after a long time) to work on the integration in EPICS of the cryoplant (this included installation of a new PLC in the cubicles, connection to the local net and debugging).

Control system.

The Personal Protection system (PPS) annual maintenance and tests were carried out several weeks in September and are completed.

In the current Integrated LIPAc schedule, it is planned to resume beam operation early November.

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Weekly LIPAc highlights (26/July-30/July)

Injector

Beam extraction with H+ at 50 keV, 20 mA for Phase B+ H+ commissioning was completed on 21 July 2021. D+ plasma conditioning was conducted from 22 to 25 July and D+ beam extraction at 100 keV, 30 mA was started on 26 July. D+ beam injection into RFQ started on 27 July.

Radio Frequency (RF) System and RFQ cavity

A short maintenance was performed for solving the issues on driver tetrode anode PS in chain 4B on 26 July (completed on 27 July morning). The RF system is operated stably at 1 Hz repetition (for 1 ms pulse). The RFQ conditioning towards 10% duty cycle has been put on-hold to prioritize beam operation.

Beam Instrumentation

A remote session with Ciemat was conducted on 26 July to improve the HEBT SEM Grid control. BPMs were not available at the beginning of the operations: they are progressively back on line and further checks are ongoing (it is necessary to debug a recent software update).

Control System

The Cryoplant LCS (Local Control System, stand-alone system at present) is under integration into LIPAc control system. We were happy to welcome after a long time an EU expert that will support this activity on-site in July and August.

LIPAc beam operation: H+ and D+ beam operation

The first H+ beam campaign at low current and low duty cycle has been successfully completed end of last week, fulfilling all the objectives defined in the Experimental Program, namely:

  • Injector test @20 mA extraction current (~10 mA at the exit of the LEBT)
  • Chopper test (100 msec)
  • LEBT (Low Energy Beam Transport) solenoid optimization (to maximize RFQ ACCT – AC Current Transformer- reading)
  • Transport up to HEBT FC (High Energy Beam Transport Faraday Cup)
  • Transport up to Beam Dump (monitor ACCTs) (~8 mA at the entrance)
  • Optimization of the dipole field
  • Beam dimension (by DPlate slits and HEBT SEM grid) and profile measurement (by HEBT SEM grids)
  • Alignment checks (screening with quadrupoles magnets)
  • Diagnostic checks

The D+ transport started beginning of this week. D+ extraction and transport up to the Beam Dump was successful with a good quality of the beam and excellent fraction of transported current (almost 80% from LEBT ACCT up to HEBT ACCT).

The following experimental program is being followed and target to be completed end of July:

  • Injector test @30 mA extraction current (~20 mA at the exit of the LEBT)
  • Chopper test (60 and 100 msec)
  • Source current adjustment
  • Transport up to HEBT FC
  • Transport up to Beam Dump (~16 mA at the entrance)
  • Optimization of the dipole field
  • Beam dimension (by DPlate slits and HEBT SEM grid) and profile measurement (by HEBT SEM grids)
  • Alignment checks (screening with quadrupoles magnets)
  • Buncher 2 phasing for BPM measurements and checks
  • Diagnostic checks as much as possible
  • Check of the beam barycenter and tuning of steerers

Next week the summer maintenance period will start. LIPAc operation is scheduled to resume end of September.

H+ beam transport along the LIPAc and beam current measurements along the line
D+ beam transport along the LIPAc and beam profile measurement at the SEM grid.
D+ beam transport along the LIPAc and beam current measurements along the line

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Weekly LIPAc highlights (12/July-16/July)

Injector

Beam extraction with H+ at 30 mA / 50keV was performed and the reproducibility of the beam characteristics was confirmed. Afterward the beam extraction with D+ was started and a current of 30 mA was achieved. The phase B+ commissioning can be started with such beam conditions.

Radio Frequency (RF) System and RFQ cavity

On HVPS 4B (High Voltage Power Supply) an interlock for Final Anode Voltage minimum is triggered when cavity voltage is set at D+ level and duty cycle is higher than 5%. With cavity voltage at H+ level the RF injection is stable. For resuming beam operation the RF system was initially set to inject RF power at H+ level in the RFQ according to the experimental programme.

Next Monday a number of interventions is scheduled on HVPS 4B to solve the issue.

Medium Energy Beam Transport (MEBT)

Further investigation of buncher #2 voltage instability at high voltages continues. Buncher#2 is not planned to be used at the beginning of phase B+ campaign.

Functional tests of updates of the LLRF (Low Level RF) SW for the timing signal have been completed.

High Energy Beam Transport (HEBT)/Beam Dump

A number of additional safety measures for HEBT vacuum system was implemented through remote session with CIEMAT. A final remote session with CIEMAT was held on July 13 for debugging dipole magnet PLC, HEBT vacuum OPI and BD instrumentation OPI.

Beam Instrumentation

The preparation of the HW for monitoring the HEBT FCU (Faraday Cup) during Phase B+ was completed. The HEBT FCU is placed before the dipole magnet (bending point of the line).

Control System

MPS (Machine Protection System) complete checks in preparation of beam operation were carried out smoothly on July 13 and 14. These tests needed the HEBT/BD LCS (Local Control System) to be connected to the LIPAcnet and all HEBT/BD OPI migrated into the integrated OPI.

Start of beam operation !

On Thursday 15 July a first dry rehearsal of the operational procedure was carried out starting at 13h30, by more than 10 operators distributed over two control rooms and with remote participation from Europe and offices on-site.

Afterwards the actual beam injection started and a first H+ beam was transported successfully through the accelerator line up to the HEBT FCU, where the oscilloscope signal confirmed the impact of the beam.

According to the experimental programme the beam current ratio for H+ and D+ should be 1:2 and the selected ion source parameters were: @20 mA injector extracted current for 10 mA H+ current at the RFQ exit and 100 us chopped pulse length with 1 Hz.

In the next days the H+ beam will be characterized at the HEBT FCU before extracting the FCU and injecting the beam all the way down to the Beam Dump. D+ beam will follow.

Integrated OPI of the LIPAc accelerator and first H+ signal detected by ACCT placed along the line.
Beam Transport line OPI and first beam detected at MEBT and DPlate ACCTs
First shot from CCR

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Weekly LIPAc highlights (28/June-2/July)

Injector

Plasma conditioning with H2 gas is regularly on-going while the vault is closed, with no issues.

The cooling panels of the LET3 and LET4 cubicles were delivered on June 28. The intervention to replace them was completed on July 01 resuming the skid operation and confirming that the solenoid PS can be turned on.

PPS (Personal Protection System) test for Injector was performed on July 02 and Injector conditioning with proton beam extraction resumed smoothly afterwards.

Radio Frequency (RF) System and RFQ cavity

After completing the cleaning and the reassembly of the flow regulators on the manifolds of the HVPS (High Voltage Power Supplies) the RF injection with RFQ condition was resumed.

RF system shows a stable behaviour and conditioning could be resumed from 5%DC and quickly ramped up to 6,7%DC (1ms pulse for 15ms period). We experienced interruptions due again to low flow rate in HVPS: when inspected the flow regulators are found to be clean, but the intervention might have jeopardized their functionality: a solution was found to modify the flow regulators and set the flow rate to higher values. The solution was successfully applied to several HVPS and will be extended to other modules whenever RF injection is interrupted by interlocks due to low mass flow.

In preparation of the beam operation validation of the upgrade of the timing system and validation of new software for the circulators are also being finalized.

Medium Energy Beam Transport (MEBT)

The conditioning of the bunchers resumed compatibly with the concurrent intervention on the RF water cooling lines: the conditioning of both bunchers was completed up to nominal voltage of 350kV. High voltage instabilities in the buncher#2 are being investigated.

High Energy Beam Transport (HEBT)

The final debugging of the HEBT/BD LCS (Local Control system) and the MPS (Machine Protection System) checks were performed with constant support by remote sessions of CIEMAT colleagues. The connection of the HEBT/BD LCS to the LIPAc-net is planned early next week and remote sessions with CIEMAT will be put on-hold until the summer maintenance period. The finalization of the HEBT vacuum machine state (not mandatory for beam operation) will be completed in September.

Beam Dump

The water circulation including the Beam Dump cartridge in its main operational mode – during beam operation – is stable. A new cooling chiller for BD skid oxygen control line has been installed and its powering was completed on July 01.

Some final adjustement on the fixing screws of the BD shield were carried out in the vault and now the installation is definitely completed.

Beam Dump shield installation in the vault completed

Beam Instrumentation

The conditioning of RGBLM (Residual Gas Bunch Length Monitor) for achieving the maximum high voltage at bottom plate electrode is completed.

Control System

MPS pre-check for the beam operation were conducted on 28 June. Only few points were found to be monitored to confirm their functionality but they are not critical for starting the initial beam operation.

The CCR OPI improvement is mostly completed and acceptance check is planned after HEBT LCS is integrated (by connecting it to the LIPAc-net).

Ancillaries

Ion exchange resins in RFQ skid water treatment line have been replaced on July 01.

MSteams operation rehearsal from CCR (Central Control Room), LCR (Local Control Room) and remote connection from Garching

A beam operation rehearsal with MS Teams was carried out on 30 June: the aim was to check the arrangement of the different operators in CCR and LCR, and perform a test of communication with Teams. All the communication was strictly conducted through Teams and direct verbal communication was avoided.

During the rehearsal the vault was closed and the RFQ conditioning started with RF injection in the cavity. F4E colleagues from Garching were also actively involved.

Operation rehearsal in CCR (Central Control Room)
Operation rehearsal in LCR (Local Control Room)
Operation rehearsal with remote participants
Operation rehearsal with remote participants and the accelerator OPI for Phase B+

Steady progress towards the beam operation!

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Weekly LIPAc highlights (14/June-18/June)

Injector

Plasma conditioning with H2 gas is regularly on-going while the vault is closed, with no issues.

Radio Frequency (RF) System and RFQ cavity

RF injection in the RFQ cavity and conditioning up to 10%DC is on-hold. It was identified that the low water mass flow rate in the HVPS (High Voltage Power Supplies) lines was caused by dirty diaphragms inside the flow regulator. The manifolds of the HVPS are being dismounted one by one and all the flow regulators are disassembled and carefully cleaned. The operation is almost completed: after reassembly on the HVPS a static pressure test and a circulation test are performed. All the HVPS cleaned so far confirmed that the nominal water mass flow rate is recovered. Actions are on-going in parallel to analyse the composition of the metallic residuals found to have clogged the flow regulators and to procure spare flow regulators as well as spare set of manifold + flow regulators. Cleaning by disassembly of flow regulators is a complex operation with high risk of breaking small parts: purchasing spares and even spare manifold is considered an important remedial action.

RF system is planned to be back operation during next week.

Medium Energy Beam Transport (MEBT)

The conditioning of the bunchers resumed compatibly with the concurrent intervention on the RF water cooling lines: the condition of buncher#1 progressed very quickly up to nominal cavity voltage (350 kV in CW for more than one hour).

High Energy Beam Transport (HEBT)

The debugging of the control system of the HEBT magnets PS has been completed with continuous support by CIEMAT with almost daily remote sessions.

The checkout of PPS signal related to the Lead Shutter was successfully completed.

Beam Dump

The water circulation including the Beam Dump cartridge in Mode 1 (main operational mode during beam operation) is stable including circulation in the BD cartiridge.

A massive operation is being carried out in the vault this Thursday and Friday with the reisntallation of the large radiation shield on the BD.

Beam Instrumentation

The stability of the accelerometers in the BD have been checked and confirmed in several remote sessions with CIEMAT.

Protection cover for RGBLM (Residual Gas Bunch Length Monitor) HV terminals have been installed in the vault and the conditioning for achieving the maximum voltage is on-going.

Control System

In view of the upcoming beam operation, the update of the required alarms to be implemented in the existing systems (BUZZER and BEAST) is on-going.

Ancillaries

The regeneration of RFQ cryopumps (4 units out of  5) and HEBT cryopumps (pumps # 1, 2 and 3) was completed smoothly.

Installation of the Beam Dump shield in the vault

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Weekly LIPAc highlights (31/May-4/June)

Radio Frequency (RF) System and RFQ cavity

RF injection in the RFQ cavity up to 5% Duty Cycle (DC) was completed and conditioning up to 10%DC has now started with remote assistance by INFN. The conditioning reached stable 6,7%DC. Wednesday a low water coolant flow rate in HVPS (High Voltage Power supply) 3A appeared again: it was not possible to adjust the mass flow rate between the main leg and the HVPS leg of the RF WCS (Water Colling System) and reach a stable regime for the WCS, and the operation was stopped. Impurities have been found in the filters at the entrance and the exit of the HVPS. An investigation on the root cause is on-going as well as an accurate cleaning. The team is working hard to resume RF-RFQ conditioning ASAP.

Medium Energy Beam Transport (MEBT)

The conditioning of the bunchers started. CW RF power was injected into the buncher #1, by following the buncher conditioning procedure and a cavity voltage of 105 kV (nominal is 350kV) was achieved.

Beam Dump

The water circulation including the Beam Dump cartridge in Mode 1 (main operational mode during beam operation) was stable. The water treatment lines are running: on the one hand the decrease of O2 is monitored, on the other hand for the first time under Mode 1 operation the pH increase was succesfully achieved by chemical injection.

During the forced stop of RF-RFQ conditioning the vault could be open and the final alignement survey of the BD cartirdge was performed, confirming that the position is within the prescribed tolerances.

Control System

The installation activities of the MPS (Machine Protection system) cables for the common chiller were completed.

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Weekly LIPAc highlights (24/May-28/May)

Injector

The injector conditioning using 6mm Plasma Electrode is currently on-hold waiting for the procurement of the cooling plates of the Power Supplies of the two Solenoid magnets. The plasma conditioning is on-going during the day when the vault is closed, and every night and weekend.

Radio Frequency (RF) System and RFQ cavity

RF injection in the RFQ cavity up to 5% Duty Cycle (DC) was completed and conditioning up to 10%DC has now started with remote assistance by INFN. The RF system is stable and no major event are observed. In the evening the system is let for night conditioning to NAT operators, while during the day time the on-site experts move to higher DC following the procedure defined with INFN requiring always tuning/adjustment of the system: for example the fine tuning of the mass flow rate balance among HVPS (High Voltage Power Supplies) and RF modules was necessary one day this week.

Medium Energy Beam Transport (MEBT)

The Personal Protection System (PPS) interface tests with the bunchers SSPA (Solid State Power Amplifiers) were successfully completed and the RF condition of the buncher cavities was started this week and is currently on-going.

Beam Dump

After the completion of the connection of the cooling loop to BD cartridge in the vault, the cartridge was successfully filled with water and the BD skid was successfully run in the two operational modes to be used for beam operation. Achieving the water circulation in the BD cartridge is a major achievement for the project and for the accelerator community worldwide! 

This could be achieved only thanks to the strenuous support of CIEMAT experts connected remote daily with Rokkasho starting from 7 am CEST and the on-site team providing an extensive support to CIEMAT experts allowing a wide amplitude of work although the time zone difference between Europe and Japan!

Rear side of the Beam Dump in the vault and its connections to the BD cooling system.

OPI of the BD cooling skid. It can be noticed that:

  • The main pump of the cooling skid BWP01 is running at nominal speed (60Hz achieving 108m3/h
  • The pump of the water treatement line (ph corection) BWP04 is operational, as well as the deareation loop (BAC01 vacuum pump).

Control System

The integration in EPICS of the MPS (Machine Protection System) signal from the Common Chiller is on-going. The enhancement of the OPI from the new CCR (Central Control Room) is showing good progresses.

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc.